# Complements of discriminants of real parabolic function singularities. II

### v3

@@ 5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well. However, over the discriminant sets of parameters, the level sets degenerate and the wave function may become irregular. The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior primarily depends on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. These homology classes depend on the connected components of the wavefront complement from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [22]).
Versal deformations of stably equivalent function singularities can be naturally identified with each other. This identification establishes a one-to-one correspondence between the components of their discriminant complements. The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number of variables, and on the parities of the inertia indices of the quadratic part of the singularity. For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas. On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or a stably equivalent function in n+2k1n+2k-1 variables generally has a different set.
@@
Almost all of these local lacunas were described in [23], and it was conjectured that there are no additional local lacunas. The above explicit enumeration of connected components of the discriminant complements proves this conjecture for all singularity classes except for $P_{8}^{2}$. For $P_{8}^{2}$, we find one new local lacuna, thus replacing the number 2 conjectured in [23] by 3.

### v4

@@ 5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals of special differential forms along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well. However, over the discriminant sets of parameters, the level sets degenerate and the wave function may become irregular. The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior primarily depends on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. These homology classes depend on the connected components of the wavefront complement from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [23], Chapter V).
Versal deformations of stably equivalent function singularities can be naturally identified with each other. This identification establishes a one-to-one correspondence between the components of the discriminant complements of these deformations. The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number $N$ of variables, and on the parities of the positive inertia indices $i_{+}$ of the quadratic part of the singularity. For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas. On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or a stably equivalent function in n+2k1n+2k-1 variables generally has a different set.
@@
Almost all of these local lacunas were described in [24], and it was conjectured that there are no additional local lacunas. The above explicit enumeration of connected components of the discriminant complements proves this conjecture for all singularity classes except for $P_{8}^{2}$. For $P_{8}^{2}$, it gives us one new local lacuna, thus replacing the number 2 conjectured in [24] by 3.